1. A statuary hall is elliptical in shape. It measures 42 feet wide and 96 feet long. If a person is standing at one focus, her whisper can be heard by a person standing at the other focus. How far apart are the two people?


2. A conic section is given by the equation 57x^2+14√3 xy+43y^2-576=0, identity and sketch the conics.

3. Find the area of the region inside both the rose curve r=sin⁡〖(2θ)〗 and the circle r=cos⁡θ. Choose the most appropriate graph and intervals to the best of your interest to solve this question

4. Sketch and identify the curve defined by the parametric equations

y=t+1, x=t^2-2t , -4≤t≤6

5. Identify and sketch the curve 9x^2-4y^2-72x+8y+176=0 using the appropriate graph. Find the equations of the asymptotes.
1
Expert's answer
2020-07-12T17:50:38-0400

1.The equation of ellipse:

"\frac{x^2}{a^2}+\frac{y^2}{b^2}=1"

For the given ellipse:

"a=96\/2=48, b=42\/2=21"

The distance between two people:

"L=2c" , where

"c=\sqrt{a^2-b^2}"

Answer:

"L=2\sqrt{48^2-21^2}=86.32" feet


2.

"57x^2+14\u221a3 xy+43y^2-576=0"

We have:

"Ax^2+2Bxy+Cy^2+2Dx+2Ey+F=0"

Canonical equation:

"A_1x^2+C_1y^2+F_1=0"

Let:

"S=A+C"

"\delta=\begin{vmatrix}
 A & B \\
 B & C
\end{vmatrix}"

"\Delta=\begin{vmatrix}
 A & B &D\\
 B & C&E\\
 D&E&F
\end{vmatrix}"

We have:

"A=57,B=7\sqrt{3},C=43,F=-576"

Then:

"S=100,\delta=2304,\Delta=-1327104"

Next, we solve the system:

"A_1+C_1=S"

"A_1C_1=\delta"

"A_1C_1F_1=\Delta"

So, we get:

"A_1=36,C_1=64,F_1=-576"

So the canonical equation:

"36x^2+64y^2=576"

"\frac{x^2}{4^2}+\frac{y^2}{3^2}=1"

Another version of equation is:

"\frac{x^2}{3^2}+\frac{y^2}{4^2}=1"

It is an ellipse with semiaxes 4 and 3.

The center can be calculated from system:

"Ax_0+By_0+D=0"

"Bx_0+Cy_0+E=0"


"57x_0+7\sqrt{3}+0=0"

"7\sqrt{3}x_0+43y_0+0=0"

"x_0=0,y_0=0"


The angle of shift of the axes is:

"tan\alpha=\frac{A_1-A}{B}=-\sqrt{3},\alpha=-60\degree"




3.Graphs intersection:

"sin2\theta=cos\theta"

"2sin\theta=1"

"\theta=\pi/6"

Then:

"S=\int_0^{\pi/6}sin^22\theta d\theta+\int_{\pi/6}^{\pi/2}cos^2\theta d\theta="

"=(\theta/2-sin(4\theta)/8|_0^{\pi/6}+(\theta/2+sin(2\theta)/4)|_{\pi/6}^{\pi/2}="

"=\pi/12-sin(2\pi/3)/8+\pi/4-\pi/12-sin(\pi/3)/4="

"=\pi/4-\sqrt{3}/16-\sqrt{3}/8"



4.

"t=y-1"

"x=(y-1)^2-2(y-1)"

"x=y^2-4y+3"

"-3\leq y\leq7"

This is parabola.



5.

"9x^2-4y^2-72x+8y+176=0"

"9(x^2-8x+16)-4(y^2-2y+1)+36=0"

"\frac{(y-1)^2}{9}-\frac{(x-4)^2}{4}=1"

This is hyperbola.

For the asymptotes:

"9(x-4)^2=4(y-1)^2-36"

"x-4=\pm\frac{2}{3}\sqrt{(y-1)^2-9}"

For "y\to\infin" :

"x-4=\pm\frac{2}{3}(y-1)"